# Quantum Takes Flight: Two-Stage Resilient Topology Optimization for UAV Networks

### v1

@@ I Introduction
To address these challenges, we propose a two-stage quantum-assisted framework that decouples computationally intensive topology optimization from online decision making. In the offline phase, we formulate the UAV topology control problem as a Quadratic Unconstrained Binary Optimization (QUBO) model, which serves as the native input format for quantum annealing (QA). This formulation allows the optimization problem to be directly mapped onto quantum hardware, enabling parallel exploration of complex solution landscapes [8]. Unlike classical methods such as simulated annealing (SA), which explore candidate configurations sequentially, QA can explore many possible topologies in parallel. Through quantum tunnelling, QA can escape from locally optimal but fragile structures, such as centralized topologies [1, 4]. Our QUBO model incorporates penalty terms to ensure load balancing, reduce node overload risk, and avoid single-point failures. In dynamic UAV networks, such imbalances and structural fragility can quickly amplify as the topology changes, leading to severe communication degradation [6]. To mitigate these effects, we employ QA with an iterative similarity penalty mechanism to generate a structurally diverse and complementary set of candidate topologies. This diversity provides the online stage with a richer set of options, increasing the likelihood of selecting a topology that best matches the current network conditions. During deployment, a lightweight classical evaluation mechanism rapidly selects the most suitable topology based on real-time link stability and Signal-to-Interference-plus-Noise Ratio (SINR), enabling fast adaptation while maintaining network quality.
@@
2) Quantum Model: The offline stage formulates the topology control task as a QUBO problem solved on a QA, which exploits quantum parallelism to explore multiple topologies simultaneously and generate a structurally diverse set of high-quality candidate topologies.
@@ II Related Work
Existing approaches either focus on static optimization or rely on heuristic and learning-based methods whose computation time is unpredictable, making them unsuitable for time-critical UAV communication. Hence, a novel framework that can efficiently generate high-quality candidate topologies offline and support rapid adaptation during flight within the operational limits of UAV hardware is needed.
@@ III System Model and Proposed Framework
$C_{ij}$ Channel capacity between UAV $i$ and $j$, calculated from free-space path loss and the Shannon model [7]
$T(\mathbf{x})$ Total achievable throughput of topology $\mathbf{x}$
@@ III-A Network Model and Objective Function
$\min_{\mathbf{x}}C(\mathbf{x})=-\alpha\,T(\mathbf{x})+\beta\,F(\mathbf{x}).$ (1)
where $T(\mathbf{x})$ represents the total network throughput and $F(\mathbf{x})$ penalizes topological fragility. The parameters $\alpha$ and $\beta$ control the trade-off between these objectives. These weights are mission dependent and can be tuned according to application priorities. For example, missions requiring continuous data transmission such as UAV monitoring favor larger $\alpha$, while long-endurance or patrol tasks prefer higher $\beta$ to emphasize structural robustness. The throughput term is defined as
@@ III-C Stage 2: Real-time Evaluation and Selection
$S_{i}(t)=w_{perf}\sum_{(j,k)\in\mathbf{X}_{i}}\log(1+\text{SINR}_{jk}(t))+w_{life}\min_{l\in\mathcal{V}(\mathbf{X}_{i})}E_{l}(t),$ (6)
where the first term reflects short-term communication performance and the second term represents network endurance. After computing the utility for all candidates, the topology with the highest score is selected:
@@
If the selected topology $\mathbf{X}_{\mathrm{sel}}$ differs from the currently deployed one (lines 11–14), a reconfiguration command is broadcast to all UAVs to update their active links accordingly. This evaluate-and-switch strategy allows the swarm to react promptly to variations in interference, mobility, or energy depletion without solving a new optimization problem in flight. The loop continues until a termination condition is satisfied, such as mission completion or network disconnection.
@@ IV Result Analysis
We evaluate the proposed two-stage framework under a controlled and fully reproducible setup. Unless otherwise stated, all results are averaged over $20$ independent initial deployments and mobility traces per setting. The simulation horizon is 30 sec with a 1 sec time step. Classical optimization is performed using D-Wave’s classical SA solver on an Apple M4 Pro CPU, and quantum solutions are obtained from Qboson qauntum annealer CPQC-550 configured with a fixed anneal schedule for fair comparison across runs [9]. Throughout this section, we use Stage 1 (offline) for candidate set generation and Stage 2 (online) for real-time selection. We compare two strategies, adopting a single optimal method as the baseline. This method computes the topology at $t=0$ using a D-Wave’s SA solver that maximizes total throughput without considering network fragility or load imbalance.
@@ IV-D Dynamic Performance Evaluation
We compare the Single Optimization Model with the proposed Dynamic Two-Stage Framework over a 30 sec mobility window with a 1 sec step. A standard air-to-air channel with standard shadowing is assumed, and link capacities are mapped to network throughput under time-division multiple access scheduling. To reflect realistic UAV operations, 3 disturbution events are injected at approximately 12 sec, 19 sec, and 25 sec, modeled as brief link outages caused by maneuvering or blockage. The single optimization model computes one fixed topology at $t=0$ sec by maximizing total throughput. The dynamic two-stage framework prepares a portfolio of 10 offline candidate topologies and performs lightweight periodic selection with hysteresis and a minimum dwell time, while modeling a short outage for topology switching. The primary metric is Performance Retention (PR), defined as $\mathrm{PR}=\frac{1}{T}\sum_{t=1}^{T}\frac{\mathrm{thr}(t)}{\mathrm{thr}(0)}$, the time average of throughput ($\mathrm{thr}$) normalized to each method’s own value at $t=0$ sec. As shown in Fig. 5, the horizontal (gray) dashed line represents the normalization reference and the shaded bands denote 95% confidence intervals over 20 runs. The dynamic two-stage framework attains PR =0.920=0.920, outperforming the Single Optimization Model at PR =0.863=0.863, which corresponds to a relative improvement of about $6.6\%$, likely to have a greater cumulative improvement on large scale settings. It also recovers faster after mission disturbances and exhibits lower variability, confirming the robustness gained from offline structural diversity combined with online selection.

### v2

@@ I Introduction
To address these challenges, we propose a two-stage quantum-assisted framework that decouples computationally intensive topology optimization from online decision making. In the offline phase, we formulate the UAV topology control problem as a Quadratic Unconstrained Binary Optimization (QUBO) model, which serves as the native input format for quantum annealing (QA). This formulation allows the optimization problem to be directly mapped onto quantum hardware, enabling parallel exploration of complex solution landscapes [8]. Unlike classical methods such as simulated annealing (SA), which explore candidate configurations sequentially, QA can explore many possible topologies in parallel. Through quantum-parallel exploration, QA can escape from locally optimal but fragile structures, such as centralized topologies [1, 4]. Our QUBO model incorporates penalty terms to ensure load balancing, reduce node overload risk, and avoid single-point failures. In dynamic UAV networks, such imbalances and structural fragility can quickly amplify as the topology changes, leading to severe communication degradation [7]. To mitigate these effects, we employ QA with an iterative similarity penalty mechanism to generate a structurally diverse and complementary set of candidate topologies. This diversity provides the online stage with a richer set of options, increasing the likelihood of selecting a topology that best matches the current network conditions. During deployment, a lightweight classical evaluation mechanism rapidly selects the most suitable topology based on real-time link stability and Signal-to-Interference-plus-Noise Ratio (SINR), enabling fast adaptation while maintaining network quality.
@@
2) Quantum Model: The offline stage formulates the topology control task as a QUBO problem solved in a QA, which exploits quantum parallelism to explore multiple topologies simultaneously and generate a structurally diverse set of high-quality candidate topologies.
@@ II Related Work
Existing approaches either focus on static optimization or rely on heuristic and learning-based methods whose computation time is unpredictable, making them unsuitable for time-critical UAV communication. Hence, a novel framework is needed that can efficiently generate high-quality candidate topologies offline and support rapid adaptation during flight within the operational limits of UAV hardware.
@@ III System Model and Proposed Framework
$C_{ij}$ Channel capacity between UAV $i$ and $j$, calculated using the Shannon capacity formula with an SNR gap approximation for practical modulation schemes [6]
$T(\mathbf{x})$ Total achievable throughput of topology (Aggregate Network Capacity) $\mathbf{x}$
@@ III-A Network Model and Objective Function
$\min_{\mathbf{x}}C(\mathbf{x})=-\alpha\,T(\mathbf{x})+\beta\,F(\mathbf{x}),$ (1)
where $T(\mathbf{x})$ represents the proxy for total network throughput and $F(\mathbf{x})$ penalizes topological fragility. The parameters $\alpha$ and $\beta$ control the trade-off between these objectives. These weights are mission dependent and can be tuned according to application priorities. For example, missions requiring continuous data transmission such as UAV monitoring favor larger $\alpha$, while long-endurance or patrol tasks prefer higher $\beta$ to emphasize structural robustness. To enable real-time quantum optimization, we adopt the Aggregate Link Capacity (sum of all active link capacities) as a part of the objective function, defined as:
@@ III-C Stage 2: Real-time Evaluation and Selection
$\begin{split}S_{i}(t)=&\,w_{perf}\sum_{(j,k)\in\mathbf{X}_{i}}\log_{2}(1+\text{SINR}_{jk}(t))\\ &+w_{life}\min_{l\in\mathcal{V}(\mathbf{X}_{i})}\left(\frac{E_{l}(t)}{E_{\text{init}}}\right)\end{split}$ (6)
where the first term reflects short-term communication performance and the second term represents normalized network endurance. After computing the utility for all candidates, the topology with the highest score is selected:
@@
If the selected topology $\mathbf{X}_{\mathrm{sel}}$ differs from the currently deployed one (lines 11–14), a reconfiguration command is broadcast to all UAVs to update their active links accordingly. This evaluation and switching strategy allows the swarm to react quickly to variations in interference, mobility, or energy depletion without solving a new optimization problem in flight. The loop continues until a termination condition is satisfied, such as mission completion or network disconnection.
@@ IV Result Analysis
We evaluate the proposed two-stage framework under a controlled and fully reproducible setup. Unless otherwise stated, all results are averaged over $20$ independent initial deployments and mobility traces per setting. The simulation horizon is 30 sec with a 1 sec time step. Classical optimization is performed using D-Wave’s classical SA solver on an Apple M4 Pro CPU, and quantum solutions are obtained from Qboson Quantum-Inspired hardware CPQC-550 configured with a fixed anneal schedule for fair comparison across runs [9]. Throughout this section, we use Stage 1 (offline) for candidate set generation and Stage 2 (online) for real-time selection. We compare two strategies, adopting a single optimal method as the baseline. This method computes the topology at $t=0$ using a D-Wave’s SA solver that maximizes total throughput without considering network fragility or load imbalance.
@@ IV-D Dynamic Performance Evaluation
We compare the Single Optimization Model with the proposed Dynamic Two-Stage Framework over a 30 sec mobility window with a 1 sec step. A standard air-to-air channel with standard shadowing is assumed, and link capacities are mapped to network throughput under time-division multiple access scheduling subject to external environmental interference. To reflect realistic UAV operations, 3 disturbance events are injected at approximately 12 sec, 19 sec, and 25 sec, modeled as brief link outages caused by maneuvering or blockage. The single optimization model computes one fixed topology at $t=0$ sec by maximizing total throughput. The dynamic two-stage framework prepares a portfolio of 10 offline candidate topologies and performs lightweight periodic selection with hysteresis and a minimum dwell time, while modeling a short outage for topology switching. The primary metric is Performance Retention (PR), defined as $\mathrm{PR}=\frac{1}{T}\sum_{t=1}^{T}\frac{\mathrm{thr}(t)}{\mathrm{thr}(0)}$, the time average of throughput ($\mathrm{thr}$) normalized to each method’s own value at $t=0$ sec. As shown in Fig. 5, the horizontal (gray) dashed line represents the normalization reference and the shaded bands denote 95% confidence intervals over 20 runs. The dynamic two-stage framework attains PR =0.920=0.920, outperforming the Single Optimization Model at PR =0.863=0.863, which corresponds to a relative improvement of about $6.6\%$, likely to have a greater cumulative improvement on large scale settings. It also recovers faster after mission disturbances and exhibits lower variability, confirming the robustness gained from offline structural diversity combined with online selection.